# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v2

@@ Abstract
Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance. However, achieving differential privacy (DP) guarantees in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. We present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph and especially consider the averaging over nearest neighbors with secure summation and individual node-wise views. Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes across different scenarios. In case the gossip averaging happens via secure summation, we show that the Rényi DP parameter growth is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
@@ 1 Introduction
In random walk-based methods (rw1; rw2; rw3), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
@@
In gossip averaging-based methods (gossip2; gossip1), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
@@
DP in Centralized FL. In centralized FL (FLbook), distributed DP (UllahCKO23; google_flblog_2023; FLDP3) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (truex2019hybrid; Erlingsson2019; bell2020secure). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
@@
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (dlattack4; dlattack3; dlattack1; dlattack2). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (HMV2015; BGTT18; XZW2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (ChanSS12).
@@
To address this, recent works propose pair-wise network DP (networkdp3; networkdp2; networkdp1), where privacy loss is analyzed between pairs of nodes. These relaxations give better utility while accounting for the structure of decentralized communication. networkdp2 give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees. Recently, bellet2025unified considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (kairouz2021practical; pillutla2025correlated) which leads to considerably more accurate privacy analysis than that of networkdp2. Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
@@
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to considering the threat model of (networkdp2), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
@@
Secure Aggregation in Decentralized Learning. While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (networkdp2), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can give a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (jayaraman2018distributed; lian2018asynchronous; jeon2021privacy). Concrete protocols for decentralized secure aggregation have been developed in recent works (sabater2022gopa; dlsa1; biswas2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (nicolas2024differentially).
@@
  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
  • We show that in case secure summation is used to carry averaging over the nearest neighbors, the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (networkdp2).
@@ 2.1 Differential Privacy
The results of (balle2018gauss) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by dong2022gaussian.
Lemma 4 (dong2022gaussian).
@@ 2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
In our results, the network DP guarantees become those of a projected Gaussian mechanism of the form
@@
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$. To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$.
We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (golub2013matrix).
@@ 2.3 Decentralized Learning and Network Differential Privacy
We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an adjacency matrix
$A\in\{0,1\}^{n\times n},\qquad A_{ij}=1\;\text{iff}\;(i,j)\in E.$
We denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
@@
Local Data and Learning Objective.
The global dataset is partitioned across nodes:
@@
where $f_{i}(\theta)$ depends only on $D_{i}$. At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$. For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by bellet2025unified as described in Section 3.4.
Dataset Adjacency.
We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
@@
Mechanism and Observations.
A decentralized mechanism is a randomized mapping
$\mathcal{M}:D\mapsto\text{(all messages exchanged during the protocol)}.$
For a subset of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{A}$ is denoted
@@
Network Differential Privacy.
Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
@@
We assume honest-but-curious nodes: they follow the protocol but analyze observed messages. As in (networkdp2), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
Gossip-Based Averaging.
Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
@@
$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},\qquad\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
For the error analysis we also consider the definition of primitive gossip matrices.
Definition 2.1 (Primitive Matrix).
@@
$W^{k}>0\quad\text{(entrywise)}.$
@@ 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., antoulas2005approximation). In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
@@
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch. 4 antoulas2005approximation). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
3.2 Analysis of Non-Adaptive Compositions
We first show how to analyze DP gossip averaging with secure summation protocols for non-adaptive node-wise functions of data. To this end, without loss of generality, consider the task of globally averaging node-wise streams of data. I.e., each node $i$, $i\in[n]$, will have a stream $x_{0}^{i},x_{1}^{i},\dots$. Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round. For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}$. Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
@@ 3.2 Analysis of Non-Adaptive Compositions
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$. Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
@@
$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$ (3.2)
Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 7 directly gives the following result.
@@
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2). Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$. Let $\widetilde{x}_{T}$ be of the form (3.2) and denote $c=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{T}\end{bmatrix}^{\top}\in\mathbb{R}^{T}$. Denote
$\Delta^{T}_{j\rightarrow i}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector. Then, from the point-of-view of a node $i$, the node $j$ is tightly $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.

### v3

@@ Abstract
Achieving differential privacy (DP) guarantees in fully decentralized machine learning is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. We present a framework for DP analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph. We present an analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes. Our main contribution is showing that the DP guarantees are those of a Gaussian mechanism, where the growth of the squared sensitivity is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation. As an application of the sensitivity analysis, we show that the excess risk of decentralized private learning for strongly convex losses is asymptotically similar as in centralized private learning.
@@ 1 Introduction
In random walk-based methods (Lopes and Sayed, 2007; Johansson et al., 2009; Mao et al., 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
@@
In gossip averaging-based methods (Boyd et al., 2006; Dimakis et al., 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
@@
DP in Centralized FL. In centralized FL (Kairouz et al., 2021c), distributed DP (Ullah et al., 2023; Kairouz et al., 2021a) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al., 2019; Erlingsson et al., 2019; Bell et al., 2020). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
@@
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al., 2025; Zhu et al., 2019; Mrini et al., 2024; Pasquini et al., 2023). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al., 2015; Bellet et al., 2018a; Xu et al., 2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al., 2012).
@@
To address this, recent works propose pair-wise network DP (Cyffers et al., 2024, 2022; Cyffers and Bellet, 2022), where privacy loss is analyzed between pairs of nodes. These relaxations give better utility while accounting for the structure of decentralized communication. Cyffers et al. (2022) give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees. Recently, Bellet et al. (2025) considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025) which leads to considerably more accurate privacy analysis than that of Cyffers et al. (2022). Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
@@
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to including the threat model of (Cyffers et al., 2022), where the neighboring nodes see the plain messages sent by their neighbors, our linear systems view allows analyzing algorithms that incorporate a secure summation protocol between neighboring nodes.
@@
Secure Aggregation in Decentralized Learning. While the privacy amplification effect coming from summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (Cyffers et al., 2022), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can give a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al., 2018; Lian et al., 2018; Jeon et al., 2021). Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al., 2022; Pereira et al., 2024; Biswas et al., 2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al., 2024).
@@
  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed via the gossip matrix and the linear systems perspective we consider.
  • We show that in case of a single node view (e.g., that of secure summation based averaging protocols), the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (Cyffers et al., 2022).
  • As an application of the sensitivity analysis, we combine the privacy bounds with the utility analysis of noisy decentralized learning (Koloskova et al., 2020) and show that the excess risk of strongly convex losses is asymptotically similar to the centralized private learning.
Relation to Recent Work by (Bellet et al., 2025). Recent work by (Bellet et al., 2025) builds on the linear systems formulation of private gossip algorithms and further represents it as a matrix mechanism (Kairouz et al., 2021b; Pillutla et al., 2025). Using this formulation, they show that the privacy guarantees extend to adaptive compositions, which we also rely on in this work.
Our contributions are complementary. While (Bellet et al., 2025) focus on adaptive privacy and general sensitivity bounds via the matrix mechanism view, we work directly in the linear systems formulation to derive analytic characterizations of both privacy and convergence for decentralized learning, which require this representation and are not obtained in their analysis. Our results show that private decentralized learning can match centralized methods in both privacy and utility.
@@ 2.1 Differential Privacy
The results of (Balle and Wang, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al. (2022).
Lemma 4 (Dong et al. 2022).
@@
An alternative and often more convenient way to analyze privacy guarantees is via Rényi Differential Privacy (RDP) (Mironov, 2017), which is based on Rényi divergence. For $\alpha>1$, the Rényi divergence of order $\alpha$ between two distributions $P$ and $Q$ is defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\!\left[\left(\frac{P}{Q}\right)^{\alpha}\right].$
A mechanism $\mathcal{M}$ is said to satisfy $(\alpha,\varepsilon)$-RDP if for all neighboring datasets $D\sim D^{\prime}$, $D_{\alpha}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq\varepsilon.$ The Gaussian mechanism with $\ell_{2}$-sensitivity $\Delta$ and noise variance $\sigma^{2}$ satisfies $(\alpha,\varepsilon)$-RDP with $\varepsilon=\frac{\alpha\,\Delta^{2}}{2\sigma^{2}}.$
We next give a result for a certain projected Gaussian mechanism, where the noise vector is projected to a smaller-dimensional subspace.
@@ 2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
In our results, the network DP guarantees become those of a Gaussian mechanism of the form
@@
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$. To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$. We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan, 2013).
@@ 2.3 Decentralized Learning and Network Differential Privacy
We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an $n\times n$ boolean adjacency matrix and we denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
@@
Local Data and Learning Objective. The global dataset is partitioned across nodes:
@@
where $f_{i}(\theta)$ depends only on $D_{i}$. At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$. For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by Bellet et al. (2025) as described in Section A.
Dataset Adjacency. We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
@@
Mechanism and View. We define a decentralized mechanism as a randomized mapping
$\mathcal{M}:D\mapsto\text{(all variables exchanged during the protocol)}.$
For a set of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{M}$ from the point-of-view of $\mathcal{A}$, i.e., the part of $\mathcal{M}(D)$ seen by nodes in $\mathcal{A}$, is denoted by
@@
Network Differential Privacy. Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
@@
We assume honest-but-curious nodes: they follow the protocol but analyze observed messages. As in (Cyffers et al., 2022), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
Gossip-Based Averaging. Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
@@
$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},$
and
$\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
Primitive gossip matrices have vanishing transient components, and we focus on this class in our asymptotic analysis.
Definition 7 (Primitive Matrix).
@@
$W^{k}>0\quad\text{(entry-wise)}.$
@@
Remark 8.
A primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value strictly less than 1.
We repeatedly use the following auxiliary lemma to bound the transient part of $W$.
Lemma 9.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Then there exists $\rho\in(0,1)$ and a symmetric matrix $R$ such that
$W=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}+R,$
$\mathbf{1}_{n}^{\top}R=0$, and for all integers $t\geq 0$, $\|R^{t}\|_{2}\leq\rho^{t}.$
@@ 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas, 2005). In particular, we consider the systems without the feedthrough term in which case their dynamics are described by the equations
@@
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{T}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[T]$ (see, e.g., Ch. 4 Antoulas, 2005). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
3.2 DP Analysis of Compositions
We show how to analyze DP gossip averaging algorithms using the linear systems view. We first focus on non-adaptive compositions, where the node-wise contributions do not depend on the state variables. The connection between the linear systems view and matrix mechanisms as established by Bellet et al. (2025) then allows extending all the results to the adaptive case. To this end, without loss of generality, consider the task of globally averaging node-wise streams of data. I.e., each node $i$, $i\in[n]$, will have a stream $x_{0}^{i},x_{1}^{i},\dots$. Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round. For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}$. Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
@@ 3.2 DP Analysis of Compositions
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $c_{t}\in\mathbb{R}$. Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of the node set $\mathcal{A}\subset[n]$, where $(\varepsilon,\delta)$ is given by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
@@
$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)$
and
$H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$, (3.2)
and $S$ is the selector matrix for $\mathcal{A}$.
Remark 11.
For example, in case we consider the threat model where all the messages sent by the neighbors are visible to a given node $i$, the selector matrix $S$ can be taken as $S=S(\bar{N}_{i})$ corresponding to the view of a closed neighborhood of node $i$. In case of secure summation over the nearest neighbors, we may take $S=e_{i}^{\top}$, i.e., the view of a single node.
Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 10 gives the following general $(\varepsilon,\delta)$-DP bound the noisy gossip averaging.
@@
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Suppose also that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2) for some $\mathcal{A}\subset[n]$. Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$. Let $\widetilde{x}_{T}$ be of the form given in Eq. (3.2). Denote
$\Delta^{T}_{j\rightarrow\mathcal{A}}$ $=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}$
$=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}$
$\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector. Then, from the point-of-view of the node set $\mathcal{A}$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
As we outline in Appendix A, due to the connection between the linear systems view and the so called matrix mechanisms, as shown by Bellet et al. (2025), Theorem 12 holds also in case adaptive cade, i.e., when the node-wise contributions $x_{t}$ depend on the state variable $\theta_{t}$. Moreover, from the spherical symmetry of the Gaussian noise it follows that the results generalize to arbitrary dimensions, when the condition $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$, is replaced by 2-norm bound for the sensitivity of node $j$’s contribution. We also remark that the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ leads to a tight $(\varepsilon,\delta)$-DP bound.